Thursday, June 4, 2020

A151 and A149 CLSL Activities Blog

3 June 2020, Tuesday - DLO

750WB console removed by CMR today.  Magnet was de-energized earlier, but is still present.

20 Aug 2019, Tuesday - DLO

A brief voltage sage occurred today at 11:17 a.m. during a lightning storm.  No known disruptions were observed.

5 July 2019, Friday - DLO

A brief power outage of apparently less than a second occurred today at 12:54 p.m.  This caused a partial shutdown of the 750NB spectrometer, but the 750WB looked outwardly normal.  Later, a problem arose with the chilled water that caused the AC to shutdown, with a significant increase in humidity.  Water condensed on cool floors, and other surfaces were sticky.  We kept the 750NB on to prevent moisture condensation.  Temperatures and humidity began to return to normal at about 4 p.m.  I estimate that conditions will be back to normal about 7 p.m. tonight.  The initial power problem was claimed to be storm-related by an F&S email.

5 Jun 2019, Wednesday - DLO

A typical "sharkfin" temperature increase occurred in A151 CLSL on Tuesday, June 4, from about 11:40 a.m. to 2:50 p.m.  The room temperature increased about 8 degrees F in about 3 hr 10 min.  At that point, the temperature decreased back to normal in about 1 hour.  The temp swing was from about 69 to 77 degrees F.  This event was recorded on the temperature trace recorded here:

Paul Barton from F&S thinks this may be related to work in the CLSL basement.

9 Oct 2018, Tuesday - DLO

Update: Hole drilling is done.  Holes were drilled at the following times for checking your NMR data:

8:40 - 8:48 a.m.
9:08 - 9:12 a.m.
11:00 - 11:18 a.m.
11:24 - 11:53 a.m.
12:13 - 12:14 p.m.
12:18 - 12:22 p.m.

Upcoming Events:

Hole Drilling in A149 (see update above).  Drilling will take place Tuesday, October 9, in the south wall of A149 CLSL for new EPR Lab equipment.   Work should begin at 11 a.m. and be completed by 1 p.m.  Three events will occur of 30 min each.

Drilling was originally intended for 7 to 9 a.m. Some brief drilling was done at 8:40 a.m. to 8:48 a.m. and 9:08 a.m. to 9:12 a.m. 

Floor Waxing of A149.  Scheduled to occur on Friday, October 12.

Removal of Plastic Draping.  Scheduled for morning of Monday, October 15.  Ideally, its removal should not affect temperatures in the main lab.

Movement of EPR magnet into A149.  The existing EPR magnet will be moved into A149 on the morning of Monday, October 15.

27 July 2018, Friday - DLO

Both dry air units back to normal after filter maintenance on unit #1 by Mark Clawson.  Lab Dew Point was always normal at -101 degrees C.

25 July 2018, Wednesday - DLO

Maintenance by Mark Clawson on one dry air unit in the basement of Chem Life.  Lab Dew Point normal at -101 degrees C.  Remote dry air monitors show one unit under maintenance.

 7 May 2018, Monday - DLO

Air conditioning was off for 4 hours the morning of Monday, May 7, but not 6 a.m. to 10 a.m., as was scheduled.  Instead, the outage occurred from about 8:30 a.m. to 12:30 p.m., as the temperature data reveal. Room temperature rose to 80 degrees F.  Ken W. castigated the F&S staff for not being on schedule, and not telling anyone.

3 May 2018, Thursday - DLO

Air conditioning will be off for 4 hours the morning of Monday, May 7, 6 a.m. to 10 a.m.  Depending on weather, room temperature may rise a few degrees.

30 Apr 2018, Monday - DLO

Update (3 p.m.):  Operations back to normal in CLSLA, and NMR Lab was never affected.

Chem Life A151 not affected by events on 3rd floor involving a hood fire this afternoon.  Monitors normal for room temperature, power, dry air, and helium bag.

17Jan 2018, Wednesday - DLO

On Tuesday, January 23, 7:30 - 8:30 a.m., two planned power sags of up to 10 V, for up to 2 sec each.  Instrumentation shutdown intended.

12 Jan 2018, Saturday - DLO

Small chance of an unplanned power outage possible for Tuesday, January 16, 7 a.m. to 5 p.m.

Dry air pressure drop occurred the afternoon of Friday, January 11 between 1 p.m. and 5 p.m.  Back-up nitrogen should have covered this pressure drop, but data acquired then should be examined.

8 Jan 2018, Monday - DLO

Small chance of an unplanned power outage possible for Tuesday, January 16, 7 a.m. to 5 p.m. 

20 Oct 2017; DLO

A voltage sag occurred in A151 CLSL on Friday, October 20.  It does not appear to have affected the NMR Lab in A151 CLSL. The sag was apparently 5.1% for 50 ms.  Our on-line data indicates that it occurred today at 14:47:50.  The reason has not been determined.


7July 2017; DLO

The electrical outage conducted on the 2nd floor Thursday, July 6, resulted in an unexpected Internet outage for the NMR Lab from 6:15 a.m to 8:48 a.m., an outage that included other parts of CLSLA.  The 750NB had to be re-mounted to Samba.  Temperature data and dry air data were not acquired during this time interval.  An oscilloscope monitoring the 750WB could not be viewed over the Internet starting at 6:15 a.m.  It was not fully understood that the Internet closet on the 2nd floor would be affected by the power outage, but it clearly was.


3 July 2017; DLO

An electrical outage is planned on the 2nd floor for 6 to 10 a.m. on Thursday, July 6.  Not expected to affect the NMR Lab in A151 CLSL. 


14 June 2017; DLO


Last night, the cooling failed in A151.  Here are some key observations:

The temperature data for A151 shows a distinct increase at 10:30 p.m.  One might guess that something turned off at that time.

The temp increased from 68 F to 82 F from 10:30 p.m. to 4:30 a.m.  That’s 14 degrees F in 7 hours.

John Rosheck did us a huge favor by sending an mail to the F & S help desk at 12:05 a.m.  I am unaware of any overnight response to this that would have set the heating event to a cooling event about 4:30 a.m.

    From about 4:30 a.m. to 6:30 a.m., the temperature recovered to the original 68 degrees F.
    John noticed an event in the power monitoring for A151.  At about 10:30 p.m., a small drop in the monitored power occurred.  At about 4:30 a.m., a small increase occurred.  We don’t know that this means.  You have to look carefully, but the signature is there.  This matter is apparently resolved,  but we are not sure exactly what happened our how it could be prevented.  Also, we are not sure how it was discovered, as F&S wrote to John and said he had contacted IT, not F&S.  An F&S person named Paul Barton thought compressors went off at some point.  Times not verified.

5 June 2017; DLO

No special activities planned from this date forward.  Any changes will be posted here.